OpenStudy (anonymous):

10. The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Write the linear equation for this problem in y=mx+b form. Note: this equation has a negative slope (m). 11. The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Make a T-table for the equation for this problem, using the x-values 0, 5, 10. Type in the ordered pairs you get in the T-table as your answer.

OpenStudy (anonymous):

10. temperature (t)=56 -1h (h=hour) so t=56-h if we are drawing a graph we place t on the y axis, and h on the x axis. so we can replace t by y, and h by x giving us y=56-x gives y= -x+56 on arranging in y=mx+b form.

OpenStudy (anonymous):

note m= -1, b=+56

OpenStudy (anonymous):

& what is 11?

OpenStudy (anonymous):

just wondering how to do the graph on here. The maths of it is easy though having done number 10.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0, number), (20, number), (40, number). like this is but not these numbers

OpenStudy (anonymous):

x 0 5 10 y =56-0 56-5 56-10 so y= =56 51 46

OpenStudy (anonymous):

so ordered pairs are (0,56); (5,51); (10,46)

OpenStudy (anonymous):

& can you answer theses two? 12.The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Use the graph you made for the equation for this problem, to approximate the answer to the following question: What temperature was it 15 hours later, according to your graph? 13.The temperature was 56 degrees this morning, but dropped one degree per hour through the rest of the day. Read over the problem and look at the graph you made to decide how to answer this question. Would it make sense to predict the temperature in 45 hours? Why or why not?